OpenStudy (anonymous):

What is the simplest form of the number? 32^4/5?

OpenStudy (solomonzelman):

\[\huge\color{red}{a^{b/c}=\sqrt[c]{a^b}}\]

OpenStudy (anonymous):

You can evaluate it and get a decimal or use what @SolomonZelman just presented, correctly. "Simplest" is in eye of beholder.

OpenStudy (solomonzelman):

\[\huge\color{red}{\sqrt[5]{32^4}}\]\[\huge\color{red}{\sqrt[5]{(2^5)^4}}\]\[\huge\color{red}{\sqrt[5]{(2^4)^5}}\]\[\huge\color{red}{2^4}\]\[\huge\color{red}{16}\]

OpenStudy (solomonzelman):

See the process?

OpenStudy (anonymous):

nice!

OpenStudy (anonymous):

beautiful, I was wrong on my answer. dang

OpenStudy (solomonzelman):

ANYTIME, guys.

OpenStudy (solomonzelman):

that's all, not hard, ah?